Project

General

Profile

NDOS CommissionMtng 20120820

Attending: Leon, Jan, Denis, Zukai, Minerba, Kurt, Mat, Peter Andrew.

Requests:
Extended noise study (Zukai)
Inquired about DSO scan for multiple channels to veto cosmic (not technically feasible) instead try subtracting the cosmic from the noise
http://nova-docdb.fnal.gov:8080/cgi-bin/ShowDocument?docid=7864
We will identify a set of channels and acquire this data on Wednesday.

UPDATES:
1.  Detector Update
Last wek APD removal: 21
Test results at Caltech:
7 tested fine
7 fail breakdown
7 nuanced behave
Working to correlate with observed features next

Leon will be at Fermilab next week for installations (parylene)

2.  DataCheck Update

Jan - Working on FEB shutoff with coating.
http://nusoft.fnal.gov/nova/datacheck/febshutoff/2012-08-10_2012-08-20.txt
Last Tuesday - DCM 1-1: many shut off
Run Aug 10th - 1 GB (high data rates)

4.  Installation Plans for the week: None

5.  DAQ Update and Plans for the week
New DCM Application release tried last week: Reconfiguration issue.
New DCM firmware trouble. Internal protocol error.
ProgUtils and DCM App (rolled back) due to DSO incompatibility. Scans halt on pix 4 of scans (Denis is investigating)

No plans this week.

6.  DCS Update and Plans for the week

Gennadiy is back; email message work
Steve drier system alarm integration work .

AOB
Minerba requested to give a tour on tuesday or Wednesday